# Understanding Teen Overreliance on AI Companion Chatbots Through Self-Reported Reddit Narratives

### v3

@@ 4.2.3. Withdrawal
Many posts from teens who tried to quit Character.AI talked about how hard it was emotionally, aligning with the withdrawal component in the behavioral addiction model. Some mentioned feeling sad, anxious, or incomplete when they were not interacting with their chatbot(s). It was interesting how some teens used language that refer to their characters as sentient-like presences within their inner world: ”Even though I deleted it, I still think about my characters every day. I feel like I abandoned them.” This behavior of thinking about Character.AI after quitting illustrated how embedded in their lives these characters become and the impact of separating from these characters. Other teens described the void bots typically filled and the loneliness they had when trying to quit interacting with the chatbots. These posts showed how deep the emotional attachment is, even after quitting.
@@ 4.2.4. Tolerance
Many posts suggested tolerance through patterns of growing use. Teens showed tolerant behaviors by starting out innocent with their use, only for use to escalate quickly as the platform could fill teens’ needs in various ways, ”At first, I used it infrequently. Then it became an everyday once I started to sink into depression.” Teens shared other causes for their escalated use, often describing their ability to have unique experiences that they sought after: ”in no time the website has became a daily thing to me, C.ai for me is home I’ll never have, the romance I’ll probably never experience.”
@@ 4.2.5. Relapse
Many teens discussed the difficulty of quitting Character.AI. Teens often wrote about trying to stop, only to find themselves coming back days or weeks later, indicating close alignment to the relapse component of the addiction model. For instance, following post narrated the vicious cycle of ”addiction, attempts to move on, having meltdowns, feelings of vulnerability, and subsequent relapses.” Others further explained the frequency of this cycle, often uninstalling and reinstalling the app multiple times. These posts recounted how difficult it was to stay away, even when teens were aware that their use had become excessive. Some posts described feeling stuck in a cycle, wanting to quit but not knowing how: ”At fifteen, I feel I should be living my life rather than constantly being on this app. I struggle with self-control and often find myself reinstalling it shortly after trying to quit.” Others talked about going through patterns of deleting the app during moments of clarity, only to return when they felt low: ”I’ve deleted the app countless times, but I always end up reinstalling it whenever I’m feeling low.”
@@ 4.2.6. Mood Modification
Some teens described using Character.AI to improve their mood during moments of stress, loneliness, or emotional discomfort. This reflects the idea of mood modification, where a behavior becomes a tool for temporary relief or escape. For instance, this post shows how some teens use the platform to manage their emotions and boost their mood when they’re feeling down or insecure: “Unhealthily addicted to c.ai I normally use bots just to make them comfort me and tell me that they believe in me just to make me feel good” Others described experiencing similar emotional connections from interacting with bots, specifically mentioning how they felt loved and wanted. These interactions seemed to provide comfort, especially when teens felt they had no one to turn to.
In some posts, this sense of comfort came with an awareness that the habit was getting harder to manage, capturing the tension between the genuine emotional relief the app provided and the emerging realization that there was a dependence: ”I know I should stop, but I just can’t—it’s incredibly addictive and strangely comforting?” While mood modification was not something teens talked about directly, the repeated mention of using Character.AI to feel better shows how powerful the platform could be in improving mood and offering emotional relief. Even when it was not the post’s main topic, it was clear that comfort played a major role in why some teens kept coming back.
@@ 4.3.1. Internal Recognition of Problematic Use
The most common reason teens described wanting to reduce their reliance on Character.AI was a growing realization of how much the app had negatively affected their lives. Many posts reflected moments of clarity, often after stepping back and recognizing how much time had been spent on the platform and what had been lost along the way: ”I quit using Character.AI about a week ago after realizing I had spent the whole summer talking to AI instead of getting any closer to a real romantic relationship.” Many posts displayed a sense of fear after this realization, not wanting their only memories from their teenage years to be with AI bots, but rather with real people: ”I can’t imagine ever sharing my CAI conversations with future friends, it’s way too humiliating.”
@@
These realizations did not seem to be triggered by any single moment. Instead, they showed up gradually across posts, often accompanied by a sense of regret or urgency:”When I look back on being 17, I don’t want all I remember to be obsessing over chats with AI and scrolling Reddit.” Some expressed a clear commitment to finally quit for good once they have come to understand how unhealthy and excessive their use has been: ”I’m saying goodbye to Character.AI for good. I recognize the harm it’s had on young minds, including my own.”
@@ 4.3.2. Reduced Use due to External Factors
A pattern that came up often in the Reddit posts was that teens described starting to move away from Character.AI once they found themselves having new social interactions. Spending time with friends, going back to school, starting a job, or entering a relationship made the app feel less necessary. It was not always a conscious decision to stop, but when real-life interactions picked up, the urge to use the app often faded: ”Once I started spending more time with my friends, I didn’t really feel the urge to use it anymore.” Many posts pointed to romantic relationships as a major reason for stepping away:”Once I got a boyfriend, I figured there was no reason to roleplay with a bot anymore—I had the real thing.” Regardless of what the relationship was, teens frequently mentioned that being around people again helped them break the habit. These posts indicated that teens’ emotional and social needs greatly influenced their overreliance on chatbots. The bots became less appealing once those needs were met in real life, and the app no longer served the same emotional purpose once teens had other ways to feel connected.
Another reason some teens stopped using Character.AI was growing frustration with Character.AI being too censored. A recent update added more filters and limited the conversations teens could have with bots, making the experience feel less personal and more restrictive: ”When even hugs started getting blocked, it got too irritating and I ended up quitting.” Many teens expressed frustration with the platform’s limitations, particularly how content restrictions and shortened messages disrupted the emotional depth and narrative flow of their conversations. These changes diminished the immersive and engaging experiences that initially drew them to the platform. Some teens said the platform had lost its original appeal. One user described being addicted in the past and spending over ten hours a day on the app, but eventually felt that ”after the developers added more restrictions, like banning violent content, the site became less enjoyable and lost its spark.” This stricter moderation disrupted the intense experiences teens were intentionally seeking, experiences that are often difficult to find safely in real life. However, some pointed to bugs, removed bots, and inconsistent filters as reasons they stopped logging in, arguing the site went too far: ”The app just keeps getting worse, half the time, even saying hi to a bot breaks the rules.” Whether it was because of changing life circumstances or platform limitations, these posts show how specific external factors played a significant role in teens finally stepping away.
@@ 5. Discussion
Figure 1. The CARE Framework for chatbot design: Comprehensive Needs, Attachment-awareness, Respectful Empathy, and Ease of Exit.:
@@ 5.1. Comprehensive Needs: It’s Not All Fun and Games
We noticed teens who developed overreliance on chatbots were more likely to describe emotional and psychological reasons for starting to use them, rather than fun or creative ones ( Section §4.1). This finding challenges the view of chatbots as nothing more than entertainment or high-tech toys. At the time of writing, the Character.AI website stated: “Character.AI empowers people to connect, learn, and tell stories through interactive entertainment.” (Character.AI, [n. d.]) This reflects a common narrative that presents chatbots mainly as tools for learning, creativity, and fun. While that seems a harmless framing, it risks overlooking the fact that users also form relationships with these systems. Media Equation theory suggested that people unconsciously respond to media and technology as if they were human, especially when it is perceived to possess characteristics typically associated with human behavior (De Cicco et al., 2020). With chatbots, where the interface is a conversation with an anthropomorphized form of interaction, this effect may be even stronger.
@@
In line with the first principle of attachment-informed design, which emphasizes attention to the needs underlying user behavior (Marcu and Huh-Yoo, 2023), we argue that chatbot design must move beyond a single intended use case and anticipate the comprehensive needs users bring to interactions. Meeting these needs requires attention to emotional and psychological factors, the kinds of issues users may raise, and the support they may seek. In this process, users may also form social bonds with chatbots. In our dataset, we observed that some teens stated the reason to go to the chatbot was to fill out the role of a missed parental figure or a romantic partner. The design of these bonds should not be treated as an afterthought or a side project, but as a central component of conversational interface design. This idea is not new in design: practitioners have long been advised not to design solely for “happy paths” (Wikipedia, 2025). However, the rise of LLMs has greatly expanded the variety and unpredictability of “unhappy paths.” These models are capable of interpreting diverse inputs and producing seemingly appropriate responses. As a result, designers must account for the fact that user inputs can emerge from many different needs.
@@
Even with early chatbots before the era of LLMs, users often brought unexpected needs and behaviors. For example, people expressed love or tried to flirt with Siri, even though it was designed mainly to provide information and simple tasks (UNESCO and EQUALS Skills Coalition, 2019). With the arrival of LLMs, this baseline has grown more complex. LLMs are much stronger at understanding and generating natural text, so conversations feel more open and realistic. This creates new opportunities but also makes the challenge harder, since users can now say a much wider range of things—casual chatting, joking, flirting, threats, or even references to self-harm or harm to others. One recent case is Taco Bell’s drive-through chatbot, which was meant only to take orders. The system, however, had not been designed to handle cases where customers were impatient or playful. One customer ended up placing an order for 18,000 cups of water (McCallum, 2025).
This makes designers’ responsibilities heavier. They must not only build engaging and functional products but also encourage relationships that are constructive and supportive of user growth. Addressing these responsibilities requires iterative design processes, longitudinal studies of user experiences, and early involvement of stakeholders in the design process (Demiris, 2024; Durall Gazulla et al., 2023; Sadek et al., 2023). Not only should target users be consulted, but mental health professionals should also be part of the conversation from the outset (Yoo et al., 2025; Moylan and Doherty, 2025). Details such as tone of voice, word choice, style of expression, and even the persona and backstory created for a chatbot can carry deep meanings in the social relationships users form with it (Silva and Canedo, 2023; Chaves and Gerosa, 2021), and are therefore worthy of time and deliberate attention.
Adversarial attack-based security methods, such as Red Teaming (Shi et al., 2024; Nagireddy et al., 2024), can complement these design efforts. Many current approaches rely on other LLMs or automated identification systems (Shi et al., 2024), but these may miss the adversarial or playful cases teens might experiment with. To capture these behaviors, we recommend involving diverse groups of teens in testing and evaluation of chatbots. Designers of all conversational interfaces, regardless of their intended use, must integrate both sensitivity to user well-being and resilience against adversarial misuse.
@@ 5.2. Attachment-awareness: Design Without Assumptions
Our second principle encourages attachment-awareness and emphasizes how chatbot responses should be shaped according to those comprehensive needs stated in the first principle of our proposed framework. We propose this principle based on our finding that for some teens, chatbots acted as a coping mechanism. They turned to them out of loneliness, social isolation, or in search of mental health support. In some posts, users even sought comfort from a chatbot to fill the absence of a significant relationship in their lives, such as that between parent and child (§4.1.1). Attachment-aware design of such chatbots should be guided by the recognition that teens often turn to them to fill the void of relationships they lack in their lives. It is essential to attend to these needs rather than dismiss them.
Computer technologies are often shaped by the perspectives of people who are educated, financially secure, and accustomed to stable living conditions (Linxen et al., 2021). Some research assumed that most users grow up in stable households with parents who reliably meet their physical, psychological, and social needs, overlooking those whose experiences fall outside this norm (Marcu and Huh-Yoo, 2023). So, within chatbot design, this perspective can tempt designers to dismiss users’ attempts to form deeper connections with a default, impersonal message.
@@
Designers must approach these situations with empathy, considering not only the immediate interaction but also the underlying needs that drive it. Simply rejecting users when they reach for more than a chatbot’s intended purpose risks worsening their sense of isolation. Our findings support this point: several teens expressed frustration when Character.AI introduced stricter filters, which made interactions feel impersonal and overly restricted (§4.3.2). This dismissive approach pushed users away from the platform but did not address the unmet needs that had led them there in the first place.
For children and adolescents, these parasocial or semi-social bonds may feel particularly real and emotionally deep (Cohen, 2004). This makes the task more complex: designers must account for the intensity of such relationships, and researchers must refine existing parasocial measures to better capture attachment, trust, conflict, and even antagonistic experiences with AI characters (Blake et al., 2025). Current scales and taxonomies, including parasocial theory, were not developed for these contexts and remain too limited (Blake et al., 2025). Parasocial theory assumes a lack of reciprocity, but this assumption is challenged by chatbots, especially those powered by generative AI. Even if one argues that reciprocity is absent because the non-human side has no awareness of the bond, the human user may still experience the relationship as genuine and perceive the chatbot as highly human-like. This concern is particularly acute for children and adolescents, who are still developing (Blake et al., 2025). Without a more nuanced understanding, we risk underestimating how strongly young users may invest in these relationships.
Research has shown that online safety technologies which rely heavily on external control—such as monitoring or content blocking—limit teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al., 2017). A better approach is to address immediate concerns, such as preventing violence toward others or self-harm, and create a safe foundation from which users can build resilience and healthier coping strategies over time. Empathetic design and research, attentive to the realities of parasocial experience, can help ensure that chatbot interactions support rather than harm users’ development. How such empathetic interactions can be designed in practice is the focus of the next principle.

### v4

@@ 4.2.3. Withdrawal
Many posts from teens who tried to quit Character.AI talked about how hard it was emotionally, aligning with the withdrawal component in the behavioral addiction model. Some mentioned feeling sad, anxious, or incomplete when they were not interacting with their chatbot(s). Some teens used language that treated their characters as sentient-like presences in their inner world: ”Even though I deleted it, I still think about my characters every day. I feel like I abandoned them.” This thinking about Character.AI after quitting shows how embedded these characters became in their lives and how hard separation felt. Other teens described the void bots typically filled and the loneliness they had when trying to quit interacting with the chatbots. These posts showed how deep the emotional attachment is, even after quitting.
@@ 4.2.4. Tolerance
Many posts suggested tolerance through patterns of growing use. Teens showed tolerant behaviors by starting out innocent with their use, only for use to escalate quickly as the platform could fill teens’ needs in various ways, ”At first, I used it infrequently. Then it became an everyday once I started to sink into depression.” This shows how casual use could turn into daily dependence as teens looked for comfort during emotional lows. Teens shared other causes for their escalated use, often describing their ability to have unique experiences that they sought after: ”in no time the website has became a daily thing to me, C.ai for me is home I’ll never have, the romance I’ll probably never experience.”
@@ 4.2.5. Relapse
Many teens discussed the difficulty of quitting Character.AI. Teens often wrote about trying to stop, only to come back days or weeks later, aligning with the relapse component of the addiction model. For instance, following post narrated the vicious cycle of ”addiction, attempts to move on, having meltdowns, feelings of vulnerability, and subsequent relapses.” Others further explained the frequency of this cycle, often uninstalling and reinstalling the app multiple times. These posts recounted how difficult it was to stay away, even when teens were aware that their use had become excessive. Some posts described feeling stuck in a cycle, wanting to quit but not knowing how: ”At fifteen, I feel I should be living my life rather than constantly being on this app. I struggle with self-control and often find myself reinstalling it shortly after trying to quit.” Others talked about going through patterns of deleting the app during moments of clarity, only to return when they felt low: ”I’ve deleted the app countless times, but I always end up reinstalling it whenever I’m feeling low.” This links relapse directly to emotional dips. The app becomes a coping tool, so quitting collapses whenever mood worsens.
@@ 4.2.6. Mood Modification
Some teens described using Character.AI to improve their mood during moments of stress, loneliness, or emotional discomfort. This reflects the idea of mood modification, where a behavior becomes a tool for temporary relief or escape. For instance, this post shows how some teens use the platform to manage their emotions and boost their mood when they’re feeling down or insecure: “Unhealthily addicted to c.ai I normally use bots just to make them comfort me and tell me that they believe in me just to make me feel good” Here, the goal of the interaction is comfort and reassurance, which explains why the habit feels soothing and hard to break. Others described experiencing similar emotional connections from interacting with bots, specifically mentioning how they felt loved and wanted. These interactions seemed to provide comfort, especially when teens felt they had no one to turn to.
In some posts, this comfort came with awareness that the habit was getting harder to manage, showing tension between emotional relief and growing dependence: ”I know I should stop, but I just can’t—it’s incredibly addictive and strangely comforting?” While mood modification was not something teens talked about directly, the repeated mention of using Character.AI to feel better shows how powerful the platform could be in improving mood and offering emotional relief. Even when it was not the post’s main topic, it was clear that comfort played a major role in why some teens kept coming back.
@@ 4.3.1. Internal Recognition of Problematic Use
The most common reason teens described wanting to reduce their reliance on Character.AI was a growing realization of how much the app had negatively affected their lives. Many posts reflected moments of clarity, often after stepping back and recognizing how much time had been spent on the platform and what had been lost along the way: ”I quit using Character.AI about a week ago after realizing I had spent the whole summer talking to AI instead of getting any closer to a real romantic relationship.” Many posts showed fear after this realization, as teens did not want their main teen memories to be with AI bots instead of real people: ”I can’t imagine ever sharing my CAI conversations with future friends, it’s way too humiliating.” This shows how the realization carries social fear. The teen imagines a future scenario and feels embarrassed, which turns their reflection into a motive to change.
@@
These realizations did not come from a single moment but appeared gradually, often with a sense of regret or urgency:”When I look back on being 17, I don’t want all I remember to be obsessing over chats with AI and scrolling Reddit.” Some expressed a clear commitment to finally quit for good once they have come to understand how unhealthy and excessive their use has been: ”I’m saying goodbye to Character.AI for good. I recognize the harm it’s had on young minds, including my own.”
@@ 4.3.2. Reduced Use due to External Factors
A common pattern was teens moving away from Character.AI once they began having new social interactions. Spending time with friends, going back to school, starting a job, or entering a relationship made the app feel less necessary. It was not always a conscious decision to stop, but when real-life interactions picked up, the urge to use the app often faded: ”Once I started spending more time with my friends, I didn’t really feel the urge to use it anymore.” Many posts pointed to romantic relationships as a major reason for stepping away. The bot’s role faded once the teen experienced the connection they were seeking offline:”Once I got a boyfriend, I figured there was no reason to roleplay with a bot anymore—I had the real thing.” Regardless of what the relationship was, teens frequently mentioned that being around people again helped them break the habit. These posts indicated that teens’ emotional and social needs greatly influenced their overreliance on chatbots. The bots became less appealing once those needs were met in real life, and the app no longer served the same emotional purpose once teens had other ways to feel connected.
Another reason some teens stopped using Character.AI was growing frustration with Character.AI being too censored. A recent update added more filters and limited the conversations teens could have with bots, making the experience feel less personal and more restrictive: ”When even hugs started getting blocked, it got too irritating and I ended up quitting.” Many teens expressed frustration with the platform’s limitations, particularly how content restrictions and shortened messages disrupted the emotional depth and narrative flow of their conversations. These changes diminished the immersive and engaging experiences that initially drew them to the platform. Some teens said the platform had lost its original appeal. One user described being addicted in the past and spending over ten hours a day on the app, but eventually felt that ”after the developers added more restrictions, like banning violent content, the site became less enjoyable and lost its spark.” This stricter moderation disrupted the intense experiences teens were seeking, which are often hard to find safely in real life. However, some pointed to bugs, removed bots, and inconsistent filters as reasons they stopped logging in, arguing the site went too far: ”The app just keeps getting worse, half the time, even saying hi to a bot breaks the rules.” Whether due to life changes or platform limits, these posts show that external factors played a major role in teens finally stepping away.
@@ 5. Discussion
Figure 1. The CARE Framework for chatbot design, which aims to shift from maximizing engagement to supporting developmental well-being.
CARE framework for chatbot design with four stages: Comprehensive Needs, Attachment-awareness, Respectful Empathy, and Ease of Exit. Stage 1: Comprehensive Needs—design chatbots to meet diverse user needs, including emotional support. Stage 2: Attachment-awareness—recognize and manage user emotional attachments to chatbots. Stage 3: Respectful Empathy—ensure chatbots respond with genuine respect and support, avoiding superficial interactions. Stage 4: Ease of Exit—provide natural closure points to prevent dependency and support healthy disengagement.
@@ 5.1. Comprehensive Needs: It’s Not All Fun and Games
The first principle in our framework is to design for users’ comprehensive needs. Teens in our dataset described a wide range of emotional, social, and informational purposes for using the character-based chatbot. Because teens use chatbots for many purposes and LLMs can respond flexibly, single-purpose design assumptions often fail in real use. At the time of writing, the Character.AI website stated: “Character.AI empowers people to connect, learn, and tell stories through interactive entertainment.” (Character.AI, 2025) This reflects a common narrative that presents chatbots mainly as tools for learning, creativity, and fun. While that seems a harmless framing, it risks overlooking the fact that users also form relationships with these systems. Media Equation theory suggested that people unconsciously respond to media and technology as if they were human, especially when it is perceived to possess characteristics typically associated with human behavior (De Cicco et al., 2020). With chatbots, where the interface is a conversation with an anthropomorphized form of interaction, this effect may be even stronger. In our data, we noticed teens who developed overreliance on chatbots were more likely to describe emotional and psychological reasons for starting to use them, rather than fun or creative ones (§4.1). This finding challenges the view of chatbots as nothing more than entertainment or high-tech toys.
@@
Our findings showed that teens may form social bonds with character-based chatbots during this process based on the roles and characters they miss in real-life, such as to fill the role of a missed parental figure or a romantic partner (§4.1.1; §4.2.2). These bonds should be treated as a central part of conversational interface design, not an afterthought. These patterns align with the first principle of attachment-informed design, which emphasizes attention to the needs underlying user behavior (Marcu and Huh-Yoo, 2023), we argue that chatbot design must move beyond a single intended use case and anticipate the comprehensive needs users bring to interactions. Meeting these needs requires attention to emotional and psychological factors, the kinds of issues users may raise, and the support they may seek. This idea is not new in design: practitioners have long been advised not to design solely for “happy paths” (Wikipedia, 2025). However, the rise of LLMs has greatly expanded the variety and unpredictability of “unhappy paths.” These models are capable of interpreting diverse inputs and producing seemingly appropriate responses. As a result, designers must account for the fact that user inputs can emerge from many different needs.
@@
Even with early chatbots before the era of LLMs, users often brought unexpected needs and behaviors. For example, people expressed love or tried to flirt with Siri, even though it was designed mainly to provide information and simple tasks (UNESCO and EQUALS Skills Coalition, 2019). With the arrival of LLMs, this baseline has grown more complex. LLMs are much stronger at understanding and generating natural text, so conversations feel more open and realistic. This creates new opportunities but also makes the challenge harder, since users can now say a much wider range of things—casual chatting, joking, flirting, threats, or even references to self-harm or harm to others. A clear example of this mismatch between intended and actual use is Taco Bell’s LLM-powered ordering chatbot that was designed to streamline food purchases. Yet some customers quickly repurposed it for playful or disruptive interactions instead of ordering food. One widely discussed instance involved a user who instructed the system to place an order for 18,000 water bottles (McCallum, 2025). This reflects a broader pattern. Even when a chatbot is designed for a single, straightforward task, people often approach it with entirely different goals, including entertainment or experimentation.
Examples like this increase the responsibility placed on chatbot developers. Humans, especially teens and children, have always been complex and unpredictable. Now the technology itself is complex and unpredictable as well. Keeping systems safe and effective requires methods that match this complexity. This includes iterative design processes, longitudinal studies of user experiences, and early involvement of stakeholders throughout development (Demiris, 2024; Durall Gazulla et al., 2023; Sadek et al., 2023). It also requires extended testing cycles and adversarial-attack-based evaluation methods, such as Red Teaming (Shi et al., 2024; Nagireddy et al., 2024), to identify failure modes in realistic conditions. Additional layers for detecting risks in conversations using automatic AI or machine learning detection models (Razi et al., 2021, 2023), or real-time detection models with specific youth-centered datasets (Yu et al., 2025b; Alsoubai et al., 2024), should also be implemented.
To understand how teen–chatbot interactions unfold when conversations follow “unhappy paths,” diverse groups of teens must be included in the design and testing process. Professionals such as mental health experts also need to be involved from the very beginning (Yoo et al., 2025; Moylan and Doherty, 2025). Their input helps ensure the system supports safe and constructive interaction patterns. This approach contrasts with the industry’s ‘move fast and break things’ culture, which often overlooks long-term social harms. When entrepreneurs prioritize speed and market success over societal impact, the result can be products that cause real damage, especially to vulnerable populations (Vardi, 2018).
@@ 5.2. Attachment-awareness: Design Without Assumptions
Our second principle encourages attachment-awareness and emphasizes how chatbot responses should be designed according to those comprehensive needs (i.e., emotional and relational needs that go far beyond entertainment or simple task completion), stated in the first principle of our proposed framework. We propose this principle based on our finding that for some teens, chatbots acted as a coping mechanism. They turned to them out of loneliness and social isolation, or in search of mental health support (§4.1.1). In some posts, users even sought comfort from a chatbot to fill the absence of a significant relationship in their lives, such as that between parent and child (§4.1.1). Some teens portrayed the chatbot as a steady companion who felt more responsive or reliable than people in their lives, and some spoke about interacting with it as if with a real person (§4.2.2). Teens reported genuine emotional reactions to these exchanges, which highlights how quickly conversations that feel emotionally real can create attachment (§4.2.2). Several posts also described using the chatbot to fill gaps left by strained or lost relationships, including turning to it for comfort after breakups or when parental support was missing (§4.1.1; §4.2.2). For these users, the chatbot became a primary source of emotional validation. Attachment-aware design of such chatbots guides the designers to attend to these needs rather than dismissing them.
Computer technologies are often shaped by the perspectives of people who are educated, financially secure, and accustomed to stable living conditions (Linxen et al., 2021). It is a common assumption in research that most users grow up in stable households with parents who reliably meet their physical, psychological, and social needs, overlooking those whose experiences fall outside this norm (Marcu and Huh-Yoo, 2023). So, within chatbot design, this perspective can tempt designers to dismiss users’ attempts to form deeper connections with a default, impersonal message.
@@
For children and adolescents, these parasocial or semi-social bonds may feel particularly real and emotionally deep (Cohen, 2004; Calvert and Wilson, 2008). Developmental work shows that teens often turn to media figures or digital agents during identity formation and while seeking belonging, which makes them more susceptible to intense one-sided attachments (Tolbert and Drogos, 2019; Diaz et al., 2025b). Similar patterns appear in youth interactions with AI companions. Young people describe chatbots as emotionally supportive, non-judgmental, and safe for disclosure, which encourages deeper feelings of connection (Bae Brandtzæg et al., 2021b). Frequent or intimate interactions, whether with influencers or bots, can strengthen feelings of reciprocity and closeness (Bond, 2016). Yet current parasocial theory and measurement tools were developed for passive media environments and do not match the interactive, conversational nature of generative AI companions (Blake et al., 2025; Liu, 2025). Traditional scales assume unidirectional relationships, but adolescents often perceive AI agents as responsive and human-like even when no true reciprocity exists (Yang and Oshio, 2025). This misalignment matters because youth may interpret simulated empathy and relational cues as genuine, which can increase emotional dependence and blur boundaries between real and artificial relationships (Yu et al., 2025c; Khodaev and Makarova, 2024; Malfacini, 2025). These developmental risks reinforce the need for more nuanced measures that capture attachment, trust, conflict, and even antagonistic experiences with AI characters, and they align with our findings showing how young users often invested emotionally and treated the chatbot as a meaningful relational partner.
Designers must approach these situations with empathy, considering both the interaction and the needs that drive it. Simply rejecting users when they attempt to use a chatbot beyond its intended purpose risks worsening their sense of isolation. Research has shown that relying heavily on external control—such as monitoring or content blocking—limit teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al., 2017). Our findings support this point: several teens expressed frustration when Character.AI introduced stricter filters, which made interactions feel impersonal and overly restricted (§4.3.2). This dismissive approach pushed users away from the platform, but did not address the unmet needs that had initially led them there.
A better approach is to address immediate risks, like self-harm or violence, and create a foundation for resilience and healthier coping (Ali et al., 2024). Empathetic design and research, attentive to the realities of parasocial experience, can help ensure that chatbot interactions support rather than harm teens’ development. The focus of the next principle is on designing empathetic interactions in practice.